MCH Corp.

Income Statement

Sales

$ 500

COGS

300

Gross Profit

200

Operating Expenses

100

Depreciation

25

EBIT

125

Interest Expense

15

EBT

110

Taxes

25

NI

85

Dividends

0

ΔRE

$85

MCH Corp.

Balance Sheet

Cash and Equivalents

$50

Accounts Receivable

100

Inventory

75

Net Fixed Assets

75

Total Assets

$300

Accounts Payable

$35

Short-Term Notes Payable

5

Accruals

10

Long-Term Debt

150

Total Liabilities

$200

Total Shareholders’ Equity

$100

Liabilities and Shareholders’ Equity

$300

What is the cash ratio?

a.

33.33%

b.

22.22%

c.

6.67%

d.

16.67%

  1. What is the days' cash held?

    a.

    60.83 days

    b.

    74.34 days

    c.

    21.91 days

    d.

    36.52 days

  2. What is the net liquid balance?

    a.

    $45

    b.

    $35

    c.

    $40

    d.

    $15
Answer #1

cash ratio = cash and equivalents / total assets = 50 / 300 = 16.67%

days cash held = cash balance / (operating expenses / 365) = 50 / (100 / 365) = 182.5 days

net liquid balance = cash and equivalents - notes payable = 50 - 5 = $45